Suppose 240 square tiles, each 25 centimeters on each side, are needed to cover a rectangular floor. The floor is 5 meters long. How wide is the floor?
Answer by richwmiller(17219)

240*25=6000 centimeters^2
6000/1000=60 meters squared
60/5=12
12 meters wide